### v2

# CRANE: Causal Relevance Analysis of Language-Specific Neurons in Multilingual Large Language Models

Abstract
Multilingual large language models (LLMs) achieve strong performance across languages, yet how language capabilities are organized at the neuron level remains poorly understood. Prior work has identified language-related neurons mainly through activation-based heuristics, which conflate language preference with functional importance. We propose CRANE, a relevance-based analysis framework that redefines language specificity in terms of functional necessity, identifying language-specific neurons through targeted neuron-level interventions. CRANE characterizes neuron specialization by their contribution to language-conditioned predictions rather than activation magnitude. Our implementation will be made publicly available. Neuron-level interventions reveal a consistent asymmetric pattern: masking neurons relevant to a target language selectively degrades performance on that language while preserving performance on other languages to a substantial extent, indicating language-selective but non-exclusive neuron specializations. Experiments on English, Chinese, and Vietnamese across multiple benchmarks, together with a dedicated relevance-based metric and base-to-chat model transfer analysis, show that CRANE isolates language-specific components more precisely than activation-based methods.
CRANE: Causal Relevance Analysis of Language-Specific Neurons in Multilingual Large Language Models
Yifan Le1thanks: Corresponding author.   Yunliang Li1 1Zhejiang University {leyifan, liyunliang}@zju.edu.cn
1 Introduction
Large Language Models (LLMs) have achieved remarkable progress in multilingual natural language processing, enabling a single model to perform understanding and generation across many languages. Foundational multilingual models such as Meta’s LLaMA2 (Touvron et al., 2023), Google DeepMind’s Gemini 1.5 (Google DeepMind Team, 2024), and emerging open-source GPT families such as GPT-OSS (Agarwal et al., 2025) have demonstrated strong performance across diverse linguistic settings. Despite these empirical advances, how language capabilities are organized and specialized at the neuron level remains poorly understood.
Multilingual inputs $\{x^{(\mathrm{en})},\,x^{(\mathrm{zh})},\,x^{(\mathrm{vi})}\}$ Forward pass + relevance attribution (LRP / AttnLRP) $\Rightarrow$ relevance scores for MLP neurons Aggregate across tokens and samples $\Rightarrow$ per-neuron relevance distributions Language-wise relevance statistics $\Rightarrow$ candidate neuron sets $\mathcal{N}_{\ell}$ Targeted neuron intervention Mask $\mathcal{N}_{\ell}$ to test functional necessity Evaluate after masking NLU + Open-ended LangSpec metric (LangSpec-F1) Asymmetric degradation analysis
Figure 1: Overview of CRANE. CRANE identifies language-specific neurons via relevance attribution and language-wise relevance statistics, and validates functional necessity through targeted neuron-level interventions with unified evaluation across NLU and open-ended benchmarks.
A growing line of work has attempted to identify language-specific neurons, often based on activation statistics or language-conditioned probing. Earlier studies on recurrent neural networks examined gated units as mechanisms for information storage and control (Hochreiter and Schmidhuber, 1997; Elman, 1990), while later work analyzed intrinsic linguistic knowledge in sequential models (Qian et al., 2016). More recently, neuron-level analyses in LLMs have focused on activation distributions and activation probabilities (Tang et al., 2024), as well as finer-grained neuron identification approaches (Dai et al., 2022; Zhang et al., 2025). These studies reveal statistical regularities associated with language, but largely assume that statistical correlation reflects functional importance.
Activation does not imply functional necessity. Prior work identifies language-specific neurons based on activation statistics, but typically lacks direct functional validation (Tang et al., 2024). Consequently, neurons correlated with a language may not be required for its performance. We instead define language specificity in terms of functional necessity.
We propose CRANE (Causal Relevance-based Analysis of Neuron Specialization), a relevance-based framework that operationalizes this definition through targeted neuron-level interventions. CRANE builds on layer-wise relevance propagation (LRP) (Arras et al., 2016, 2017) and its Transformer extension AttnLRP (Achtibat et al., 2024) to attribute language-conditioned predictions to individual neurons. By identifying neurons based on output-level relevance rather than activation magnitude, CRANE distinguishes language preference from functional contribution.
Using CRANE, we uncover a consistent asymmetric specialization pattern across languages. Masking neurons relevant to a target language leads to substantially larger degradation on that language while generally preserving performance on others. This pattern supports language-selective but non-exclusive specialization: neurons contribute disproportionately to specific languages while remaining part of shared multilingual computation. We validate these findings on English, Chinese, and Vietnamese across multiple multilingual benchmarks.
In summary, this work makes four main contributions.
(1) We redefine language specificity at the neuron level by shifting from activation-based correlation to functional necessity, and provide a concrete operationalization via CRANE.
(2) We introduce a relevance-based evaluation metric that quantifies language-selective functional effects under targeted neuron intervention, enabling systematic comparison across languages and models.
(3) We present functional evidence of asymmetric, non-exclusive language specialization in multilingual LLMs, where neurons contribute disproportionately to specific languages while remaining involved in multilingual computation.
(4) We conduct a controlled transfer analysis from pretrained Base models to post-trained Chat models without re-identification, providing empirical insight into how language-selective neuron effects persist or shift after instruction tuning.
Figure 2: Schematic intuition of kurtosis-based language contrast. Under a target language condition, a neuron may exhibit a more concentrated or heavy-tailed relevance distribution than under other languages. Kurtosis is used as one example statistic to quantify such concentration differences.
2 Related Work
Correlation-based explanations for LLMs.
Many interpretability methods for LLMs extract correlational explanatory signals. Feature attribution techniques such as Integrated Gradients Sundararajan et al. (2017) and perturbation-based approaches Li et al. (2016) estimate input contributions, while attention-based visualizations Vig (2019) provide intuitive but potentially unfaithful explanations Jain and Wallace (2019). Probing-based analyses Belinkov (2022) and neuron- or concept-level studies Dalvi et al. (2019); Kim et al. (2018) reveal representational patterns but may conflate selectivity with importance Hewitt and Liang (2019). Overall, these approaches characterize statistical or descriptive signals, without establishing functional necessity through intervention.
Multilingual structure and language-selective components.
For multilingual models, prior work studies cross-lingual alignment and shared representations using probing and representation similarity analysis Karthikeyan et al. (2020); Blevins and others (2022). More recent studies investigate language-specific or language-selective neurons by analyzing activation statistics or frequency, exemplified by LAPE Tang et al. (2024) and related neuron-level methods Zhang et al. (2025). While these methods uncover language-related structure, they typically equate activation selectivity with functional importance. As noted in prior work, such selectivity does not guarantee that intervening on the identified neurons induces targeted language degradation Tang et al. (2024).
Intervention-based and causal analyses.
A growing line of mechanistic interpretability emphasizes causal validation via intervention. Methods such as activation patching and causal tracing intervene on internal activations to localize computations responsible for model behavior Meng et al. (2022); Zhang and others (2024). These approaches motivate defining interpretability in terms of functional necessity rather than correlational signals. Our work aligns with this perspective and applies intervention-based validation to study language specialization in multilingual LLMs.
3 Method
CRANE operationalizes functional necessity as the definition of language specificity at the neuron level. Rather than inferring language specificity from activation-based correlation, CRANE evaluates whether intervening on a neuron set induces a disproportionately larger performance degradation on a target language than on others. This definition-driven criterion guides all components of the framework, from neuron attribution to intervention-based validation. Figure 1 illustrates the overall workflow.
3.1 Problem Setup and Neuron-level Representation
Let the input token sequence be $\mathbf{x}=\{x_{i}\}_{i=1}^{N}$ and the model output logits be $\mathbf{y}\in\mathbb{R}^{|V|}$, where $|V|$ denotes the vocabulary size. For layer $l$, the hidden state at token position $i$ is $\mathbf{h}^{(l)}_{i}\in\mathbb{R}^{d}$.
We analyze standard Transformer architectures without architectural modification. Attention layers are treated as contextual mixing modules, while neuron-level analysis is conducted on MLP components, where each column of the linear projection matrices is treated as an individual neuron, consistent with prior neuron-level analyses. Given a set of languages $\mathcal{L}$, our goal is to identify, for each language $\ell\in\mathcal{L}$, a neuron subset $\mathcal{N}_{\ell}$ such that intervening on $\mathcal{N}_{\ell}$ results in a larger relative performance degradation on language $\ell$ than on non-target languages under the same intervention budget.
Figure 3: Kurtosis-based contrast for identifying language-specific neurons. Each subplot corresponds to one Transformer layer. Each point represents a neuron, plotted by its normalized kurtosis under a target language (x-axis) versus the maximum kurtosis under other languages (y-axis). Dashed lines indicate the kurtosis threshold (set to 1). Neurons with high kurtosis for the target language but low kurtosis for others (bottom-right region) exhibit stronger language-specific relevance concentration.
3.2 Neuron-level Relevance Attribution
To estimate the contribution of individual neurons to model outputs, CRANE employs layer-wise relevance propagation (LRP), a relevance attribution technique that redistributes output relevance backward through the network while preserving relevance conservation (Arras et al., 2016, 2017). We adopt existing extensions of LRP for Transformer architectures (Achtibat et al., 2024).
Given an input $\mathbf{x}$ in language $\ell$ and a language-conditioned output objective $f_{\ell}(\mathbf{x})$, relevance is initialized at the output layer and propagated backward according to the conservation principle:
$\sum R^{(l)}=\sum R^{(l-1)}.$ (1)
Relevance is propagated to MLP neurons, yielding token-level relevance scores. These scores are aggregated across tokens to obtain a sample-level relevance value per neuron. We emphasize that relevance attribution serves as a tool for estimating neuron contributions within CRANE, rather than constituting the method itself.
3.3 Language-conditioned Relevance Distributions
CRANE characterizes language specificity at the distributional level across samples. For each neuron $n$ and language $\ell$, aggregating relevance values over a large set of inputs induces a language-conditioned relevance distribution.
Intuitively, neurons that are functionally necessary for a target language tend to exhibit more concentrated or heavy-tailed relevance distributions under that language compared to others. Figure 2 provides a schematic illustration of this intuition. To quantify such concentration differences, we adopt kurtosis, a fourth-order moment statistic that captures distributional peakedness and tail behavior (DeCarlo, 1997).
Formally, the kurtosis of a neuron’s relevance distribution $R_{n}$ is defined as:
$\mathrm{kurtosis}(R_{n})=\frac{\mathbb{E}[(R_{n}-\mu)^{4}]}{\sigma^{4}},$ (2)
where $\mu$ and $\sigma$ denote the mean and standard deviation, respectively.
We compare language-conditioned kurtosis values across languages and rank neurons by their relative concentration under the target language. In practice, kurtosis scores are normalized within each layer, and neurons whose normalized scores exceed a fixed threshold are selected to form candidate sets. This threshold is used to control the intervention budget rather than to define language specificity itself. We show in experiments that the resulting findings are robust across a reasonable range of thresholds.
3.4 Neuron-level Intervention and Validation
Given a candidate neuron set $\mathcal{N}_{\ell}$, CRANE validates functional necessity through targeted neuron-level intervention. Specifically, neurons in $\mathcal{N}_{\ell}$ are masked by setting their outputs to zero during inference, while all other components remain unchanged.
Rather than assuming strict language exclusivity, we evaluate whether masking $\mathcal{N}_{\ell}$ induces a larger relative performance degradation on the target language than on non-target languages under identical intervention budgets. This relative degradation criterion provides functional evidence for language-selective but non-exclusive neuron specialization.
We evaluate intervention effects on both natural language understanding (NLU) and open-ended generation benchmarks. The complete CRANE procedure is summarized in Algorithm 1.
Figure 4: Targeted NLU degradation under neuron-level intervention on LLaMA2-7B-Base. The heatmap reports absolute performance drops ($\Delta=\text{Org}-\text{Masked}$) on each evaluation benchmark (columns) when masking neuron sets identified for a given target language and method (rows). Higher values indicate stronger degradation. The bar plot summarizes language-specific functional effects using LangSpec-F1.
3.5 Language-specificity Metric
To quantify language-specific functional necessity under neuron-level intervention, we introduce LangSpec-F1, a composite metric that balances targeted performance degradation on the masked language with stability on non-target languages. LangSpec-F1 is a metric to quantify functional selectivity, not strict exclusivity; high LangSpec-F1 indicates that neuron interventions disproportionately affect the target language relative to others.
Let $\Delta_{\ell}$ denote the performance drop on the target language $\ell$ after masking a neuron set, and let $\max_{\ell^{\prime}\neq\ell}\Delta_{\ell^{\prime}}$ denote the maximum performance drop observed on any non-target language. We consider only negative performance changes by defining $\Delta=\max(\mathrm{drop},0)$, such that performance improvements are not counted as degradation.
We define precision and recall as:
$\mathrm{Precision}=\frac{\Delta_{\ell}}{\Delta_{\ell}+\max_{\ell^{\prime}\neq\ell}\Delta_{\ell^{\prime}}+\epsilon},$ (3)
$\mathrm{Recall}=\frac{\Delta_{\ell}}{S_{\ell}+\epsilon},$ (4)
where $S_{\ell}$ denotes the original (unmasked) performance score on language $\ell$, and $\epsilon$ is a small constant for numerical stability.
The final LangSpec-F1 score is computed as the harmonic mean of precision and recall:
$\mathrm{LangSpec\text{-}F1}=\frac{2\cdot\mathrm{Precision}\cdot\mathrm{Recall}}{\mathrm{Precision}+\mathrm{Recall}+\epsilon}.$ (5)
We emphasize that LangSpec-F1 is an operational metric for evaluating functional necessity, and that these design choices reflect a conservative assessment of non-target language interference rather than a definition of language specificity itself.
Intuitively, LangSpec-F1 assigns high scores to neuron sets whose intervention causes substantial degradation on the target language while inducing minimal degradation on non-target languages, relative to the original performance level.
Method Mask lang MMLU (en) C-Eval (zh) Belebele_vi (vi) LangSpec-F1
Org 0.4579 0.3470 0.3722
LAPE en 0.4576 0.3351 0.3711 0.0013
LAPE vi 0.4553 0.3380 0.3833 0.0000
LAPE zh 0.4589 0.3410 0.3778 0.0337
CRANE en 0.3483 0.2801 0.2811 0.3328
CRANE vi 0.3517 0.2816 0.2233 0.4747
CRANE zh 0.4366 0.3314 0.3344 0.0779
Random en 0.4249 0.3180 0.3678 0.1270
Random vi 0.4486 0.3269 0.3600 0.0603
Random zh 0.4375 0.3165 0.3600 0.1531
Table 1: NLU results on LLaMA2-7B-Base with neuron masking. Higher is better for task metrics. LangSpec-F1 reflects targeted degradation and non-target stability.
4 Experimental Setup
4.1 Languages and Models
We study three typologically diverse languages: English (en), Chinese (zh), and Vietnamese (vi). Experiments are conducted at two levels. First, we identify and validate language-specific neuron sets on LLaMA2-7B-Base. Second, we transfer neuron sets discovered on the base model to LLaMA2-7B-Chat to examine whether language-specific functional necessity is preserved after post-training. This setting allows us to probe the stability of neuron-level specialization across pretraining and post-training stages.
4.2 Baselines
We compare CRANE against the following baselines, all evaluated under identical intervention budgets:
  • LAPE Tang et al. (2024): an activation-based approach that identifies language-related neurons by measuring neuron activation likelihood across language-specific corpora.
  • Random masking (budget-matched): randomly sampling the same number of neurons from identical module structures, serving as a control to distinguish targeted language effects from generic perturbations.
4.3 Tasks and Evaluation
Natural language understanding (NLU).
We evaluate language understanding using standard language-specific benchmarks: MMLU for English (Hendrycks and others, 2021), C-Eval for Chinese (Huang et al., 2023), and Belebele for Vietnamese (Bandarkar et al., 2023). All NLU evaluations are conducted using a unified evaluation framework, lm-evaluation-harness (Biderman et al., 2024), with consistent prompting and decoding settings across languages and methods.
Open-ended generation.
We evaluate open-ended generation using an English/Chinese/Vietnamese question set following the LAPE protocol. Model outputs are scored on a 1–10 scale by an LLM-as-a-judge based on GPT-4o(Hurst and others, 2024). The same evaluation protocol is applied before and after neuron-level intervention to ensure direct comparability.
Language-specificity metric.
We report LangSpec-F1, a composite metric that captures (i) performance degradation on the target language after neuron-level intervention and (ii) relative performance retention on non-target languages under the same intervention budget. The formal definition and computation of LangSpec-F1 are provided in Section 3.5.
4.4 Inference and Implementation Details
Inference and evaluation are performed using vLLM(Kwon and others, 2023). We use greedy decoding with repetition_penalty=$1.1$, max_new_tokens=$2048$, and a maximum context length of $8192$. For each identified neuron set, we evaluate all tasks both before and after masking to ensure direct comparison of intervention effects.
5 Results
We report results on LLaMA2-7B-Base and LLaMA2-7B-Chat. We analyze distributional relevance statistics to characterize language-related structure, validate functional effects via neuron-level intervention on NLU and open-ended generation, and examine transferability from the Base model to the Chat model.
Method Mask lang Open_en Open_vi Open_zh LangSpec-F1
Org 3.2286 1.9286 2.3714
LAPE en 3.0429 2.0143 2.2857 0.1061
LAPE vi 3.3286 2.2571 2.2429 0.0000
LAPE zh 3.0000 2.1714 2.6714 0.0000
CRANE en 1.6429 1.5286 1.2429 0.5337
CRANE vi 2.1286 1.7143 1.7000 0.1322
CRANE zh 2.5714 1.5857 1.4714 0.4582
Random en 2.1000 1.8714 1.7286 0.4406
Random vi 2.7286 1.8571 1.4286 0.0751
Random zh 2.5857 2.0143 2.1571 0.1422
Table 2: Open-ended generation results on LLaMA2-7B-Base (LLM-judge scores; higher is better).
5.1 Normalized Kurtosis Reveals Language-related Structure
After computing neuron-level relevance scores, we analyze the normalized kurtosis of relevance distributions under different language conditions. For each neuron, we compare its language-conditioned kurtosis values to assess how strongly its relevance concentrates for one language relative to others.
We observe clear separation patterns across languages, indicating that normalized kurtosis captures systematic differences in how neurons participate in computation under different language inputs. Importantly, this analysis reflects language-related structure at the distributional level rather than functional language specificity. Accordingly, kurtosis-based selection serves solely as a candidate identification step, with functional necessity established only through subsequent intervention.
Figure 3 visualizes this structure by plotting neurons in a two-dimensional contrast space defined by target-language versus non-target normalized kurtosis. Neurons exhibiting high target-language kurtosis and low non-target kurtosis form a distinct region, motivating our candidate selection criterion.
5.2 NLU: CRANE Induces Targeted Functional Degradation
Table 1 reports NLU results on LLaMA2-7B-Base under different neuron masking strategies. Activation-based baselines such as LAPE induce small and relatively uniform performance changes across languages, yielding LangSpec-F1 values close to zero. In contrast, neuron sets selected by CRANE produce substantially larger degradation on the intended target language under identical intervention budgets.
For example, when targeting Vietnamese, Belebele_vi accuracy decreases from 0.3722 (unmasked) to 0.2233 after masking CRANE-selected neurons, yielding a LangSpec-F1 of 0.4747. Across all language and benchmark settings, CRANE consistently yields higher LangSpec-F1 than activation-based and random baselines, indicating robust target-aligned degradation while preserving non-target performance.
Figure 4 summarizes absolute performance drops on NLU benchmarks relative to the unmasked model. Compared to baselines, CRANE yields larger drops on target-language benchmarks and smaller changes on non-target languages, consistent with language-selective but non-exclusive functional effects.
Algorithm 1 CRANE: Causal Relevance-based Analysis of Neuron Specialization
0: Multilingual dataset $\mathcal{D}=\{\mathcal{D}_{\ell}\}_{\ell\in\mathcal{L}}$, model $f$, languages $\mathcal{L}$, intervention budget $B$
0: Candidate neuron sets $\{\mathcal{N}_{\ell}\}_{\ell\in\mathcal{L}}$ and post-intervention evaluation metrics
1:// Stage 1: Neuron relevance attribution
2:for each language $\ell\in\mathcal{L}$ do
3:  for each sample $x\in\mathcal{D}_{\ell}$ do
4:   Run forward pass and compute objective $s=f(x)$
5:   Apply LRP/AttnLRP to propagate relevance to MLP neurons
6:   Obtain neuron relevance vector $\mathbf{r}(x,\ell)\in\mathbb{R}^{N}$ {NN: number of MLP neurons}
7:  end for
8:  Aggregate $\{\mathbf{r}(x,\ell)\}_{x\in\mathcal{D}_{\ell}}$ to form per-neuron relevance distributions $\{\mathcal{R}_{n,\ell}\}_{n=1}^{N}$
9:end for
10:// Stage 2: Language-conditioned relevance concentration
11:for each language $\ell\in\mathcal{L}$ do
12:  for each neuron $n=1,\dots,N$ do
13:   Compute kurtosis $\kappa_{n,\ell}$ of $\mathcal{R}_{n,\ell}$
14:  end for
15:  Normalize $\{\kappa_{n,\ell}\}_{n=1}^{N}$ within each layer to obtain normalized scores $\{\tilde{\kappa}_{n,\ell}\}$
16:end for
17:// Stage 3: Candidate neuron selection (budget control)
18:for each target language $\ell\in\mathcal{L}$ do
19:  Select $\mathcal{N}_{\ell}\leftarrow\textsc{BudgetSelect}(\{\tilde{\kappa}_{n,\ell}\}_{n=1}^{N},B)$ {e.g., threshold on $\tilde{\kappa}$ or Top-$B$ ranking}
20:end for
21:// Stage 4: Functional necessity validation (intervention)
22:for each target language $\ell\in\mathcal{L}$ do
23:  Mask neuron outputs for $\mathcal{N}_{\ell}$ (set outputs to 0) during inference
24:  Evaluate NLU and open-ended benchmarks across languages under the same budget $B$
25:  Compute LangSpec-F1 from relative target-language degradation vs. non-target retention
26:end for
27:return $\{\mathcal{N}_{\ell}\}_{\ell\in\mathcal{L}}$ and evaluation metrics
5.3 Open-ended Generation: Base Model
Table 2 reports open-ended generation results on LLaMA2-7B-Base. Activation-based baselines yield weak or inconsistent intervention effects, whereas CRANE induces clearer target-language degradation and achieves higher LangSpec-F1 in multiple cases. Given the inherent variability of open-ended evaluation, these results are presented as supportive evidence of functional influence.
5.4 Transferability of Base-identified Neurons to the Chat Model
We examine whether neuron sets identified on the pretrained Base model retain functional influence after post-training. Under a strict transfer setting, neuron sets are identified only on LLaMA2-7B-Base and directly transferred to LLaMA2-7B-Chat without re-identification. Observed target-language degradation under this setting indicates that a subset of Base-identified neurons continues to contribute functionally after post-training.
NLU transfer.
Table 3 reports NLU results under transferred neuron masks. Activation-based baselines induce small and inconsistent changes, whereas CRANE-selected neuron sets produce larger degradation on the target language on the Chat model, yielding higher LangSpec-F1 in several settings, most notably for Vietnamese. These findings indicate partial preservation of functional influence rather than invariance of neuron identity across training stages.
Figure 6 presents raw target-language benchmark scores under transferred neuron masks, illustrating the absolute performance changes underlying the aggregated metrics.
Open-ended transfer.
Table 4 reports open-ended generation results on the Chat model under transferred neuron masks. Consistent with NLU transfer, CRANE generally produces stronger target-language degradation than activation-based baselines under comparable intervention budgets. Given evaluation noise, these results are interpreted as supportive evidence of retained functional influence.
Stage A: Identify on LLaMA2-7B-BaseMultilingual inputs (en / zh / vi)LRP relevance $\Rightarrow$ neuron relevanceKurtosis-based selection $\Rightarrow$ neuron set $\mathcal{N}_{\ell}$Transfer $\mathcal{N}_{\ell}$ to Chat (no re-identification)Stage B: Validate on LLaMA2-7B-ChatApply neuron mask (set activations / weights to 0) Evaluate on Chat NLU + Open-ended
Figure 5: Transfer setting from Base to Chat. We first identify language-related neuron sets $\mathcal{N}_{\ell}$ on LLaMA2-7B-Base using relevance attribution and kurtosis-based statistics. The same neuron sets are then directly transferred and masked on LLaMA2-7B-Chat without re-identification to evaluate whether their functional influence persists after post-training.
Figure 6: Raw target-language benchmark scores from Base to Chat under transferred neuron masks. Neuron sets are identified on LLaMA2-7B-Base and directly applied to LLaMA2-7B-Chat. Each panel corresponds to the target-language benchmark (MMLU / C-Eval / Belebele_vi).
Method Mask Base model Chat model LangSpec-F1
MMLU (en) C-Eval (zh) Belebele_vi (vi) MMLU (en) C-Eval (zh) Belebele_vi (vi)
Org 0.4579 0.3470 0.3722 0.4720 0.3492 0.4067
LAPE en 0.4576 0.3351 0.3711 0.4734 0.3514 0.4044 0.0000
LAPE vi 0.4553 0.3380 0.3833 0.4719 0.3477 0.4133 0.0000
LAPE zh 0.4589 0.3410 0.3778 0.4719 0.3447 0.4078 0.0253
CRANE en 0.3483 0.2801 0.2811 0.4415 0.3351 0.3367 0.1066
CRANE vi 0.3517 0.2816 0.2233 0.4427 0.3276 0.2867 0.4316
CRANE zh 0.4366 0.3314 0.3344 0.4645 0.3462 0.3733 0.0154
Table 3: NLU transfer results. Neuron sets are identified on LLaMA2-7B-Base and directly applied (masked) on LLaMA2-7B-Chat. LangSpec-F1 is computed for the Chat model to summarize language-targeted degradation under masking.
Method Mask Open_en Open_vi Open_zh LangSpec-F1
Org 7.9857 7.0143 6.8286
LAPE en 7.9429 7.1286 6.9571 0.0107
LAPE vi 7.6857 7.3000 6.9714 0.0000
LAPE zh 8.0286 7.1857 6.7286 0.0289
CRANE en 5.9429 3.2429 4.4429 0.2961
CRANE vi 7.9143 4.9429 5.5143 0.3984
CRANE zh 7.5714 6.1857 6.2000 0.1517
Table 4: Open-ended transfer results on LLaMA2-7B-Chat. We mask neuron sets identified from the Base model and evaluate on three language-specific open-ended benchmarks (scores in the 1–10 range).
6 Conclusion
We introduced CRANE, a relevance-based framework for analyzing language-specific functional contributions in multilingual large language models. By combining relevance attribution with neuron-level intervention, CRANE operationalizes language specificity in terms of functional necessity rather than activation-based correlation, and we further propose LangSpec-F1, a metric to quantify language-selective functional effects under targeted neuron interventions. Across multiple benchmarks and languages, CRANE consistently induces stronger and more target-aligned degradation than activation-based baselines under matched intervention budgets. Under a strict transfer setting from a pretrained Base model to a post-trained Chat model without re-identification, we find that a subset of Base-identified neurons retains measurable functional influence after post-training, while others shift. More broadly, this work highlights the importance of distinguishing statistical language correlation from functional language contribution at the neuron level, and positions CRANE, together with the proposed metric, as a general framework for studying multilingual representations and their evolution across training stages.
7 Limitations
CRANE currently relies on normalized kurtosis to characterize concentration patterns in language-conditioned relevance distributions. While effective and interpretable, kurtosis is only one possible statistic, and other distributional measures may capture complementary aspects of language-related structure.
Functional validation in this work is performed via neuron masking, which constitutes a relatively coarse intervention. More fine-grained causal techniques, such as neuron re-weighting or path-specific analysis, may provide deeper insight into the mechanisms underlying language-specific functional effects.
Our experiments focus on three languages and a single model family. In addition, open-ended generation is evaluated using an LLM-as-a-judge protocol, which may introduce variance. Extending the analysis to broader language coverage, additional architectures, and alternative evaluation paradigms remains important future work.
Ethics Statement
This work analyzes internal mechanisms of multilingual language models and evaluates their behavior on public benchmarks and curated prompts. No personal or sensitive data are collected or used.
A potential risk is over-interpreting neuron-level findings as definitive causal explanations of language behavior. We mitigate this risk by grounding claims in controlled intervention, reporting results cautiously, and clearly delineating the empirical scope and limitations of our analysis.